# Test-Time Model Adaptation with Only Forward Passes

### v1

@@ 4.4 More Discussions
Computational Complexity Analyses. The primary computational demand of FOA stems from $K$ (the population size in CMA) forward passes. Though FOA requires more forward passes than TENT, its independence from backward passes significantly reduces memory usage and potentially boosts overall efficiency. In Table 8, our Activation Shifting achieves almost the same efficiency as NoAdapt while outperforming T3A, i.e., 59.1% vs. 56.9% accuracy. Notably, our BP-free FOA ($K\small{=}2$) matches the accuracy of BP-based TENT with lower run time and memory, and FOA ($K\small{=}6$) further surpasses BP-based CoTTA in all aspects. Moreover, even at $K\small{=}28$, FOA maintains much higher efficiency than augmentation-based methods like MEMO (Zhang et al., 2022). In FOA, $K$ is a hyperparameter that one can select different values for performance and efficiency trade-off. Here, we do not report MEMO’s performance as it is too time-consuming and has lower accuracy than TENT per Niu et al. (2023). Note that TENT updates only norm layers, making it more efficient than full-model backpropagation.
@@
Average GPU
Method BP #FP #BP Acc. ECE Time (s) Memory (MB)
@@ 5 Conclusion
In this paper, we aim to implement online test-time adaptation without the need for backpropagation and altering the model parameters. This advancement highly broadens the scope of TTA’s real-world applications, particularly in resource-limited scenarios such as smartphones and FPGAs where backpropagation is often not feasible. To this end, we propose a test-time Forward-Only Adaptation (FOA) method. In FOA, we online learn an input prompt through a covariance matrix adaptation technique, paired with a designed unsupervised fitness function to provide stable learning signals. Moreover, we devise a “back-to-source” activation shifting scheme to directly alter the activations from the out-of-distribution domain to the source in-distribution domain, further boosting the adaptation performance. Extensive experiments on four large-scale out-of-distribution benchmarks with full precision (32-bit) and quantized (8-bit/6-bit) models verify our superiority.
Impact Statements
@@ Appendix A Related Work
$\bullet$ Backpropagation (BP)-Free TTA. In the early development of BP-free TTA, attention was primarily given to adjusting batch normalization (BN) layer statistics by computing mean and variance from testing data (Nado et al., 2020; Schneider et al., 2020). This method, however, assumes multiple test samples for each prediction. To address this, later studies proposed single-sample BN adaptation techniques, such as using data augmentation (Khurana et al., 2021), mix-up training and testing statistics (Hu et al., 2021; Lim et al., 2023). In addition to BN adaptation, other methodologies have been explored and shown to be effective, such as prototype-based classifier adjustment (Iwasawa & Matsuo, 2021), predicted logits correction (Boudiaf et al., 2022), difussion-based input image adaptation (Gao et al., 2023) and test-time batch renormalization (Zhao et al., 2023). However, since BP-free TTA does not update the core model parameters, it might exhibit limited learning capabilities, often resulting in suboptimal performance when dealing with out-of-distribution testing data. Therefore, BP-based TTA emerged as an effective solution to significantly boost the out-of-distribution generalization performance.
@@
$\bullet$ Backpropagation-Based TTA. One pioneering work of BP-based TTA is known as Test-Time Training (TTT) (Sun et al., 2020). For TTT methods, they first train a source model using both supervised and self-supervised objectives, followed by adapting the model at test time with the self-supervised objective, such as rotation prediction (Sun et al., 2020), contrastive learning (Liu et al., 2021a; Bartler et al., 2022), reconstruction learning (Gandelsman et al., 2022; Deng et al., 2024). To avoid directly altering the model training process and access to source data, Fully TTA methods update any given model via unsupervised learning objectives, such as entropy minimization (Wang et al., 2021; Wen et al., 2023; Niu et al., 2023), prediction consistency maximization (Zhang et al., 2022; Fleuret et al., 2021; Zeng et al., 2023) and feature distribution alignment (Mirza et al., 2023; Lin et al., 2023).
@@
However, BP-based TTA methods typically require multiple backward propagations for each test sample, leading to computational inefficiency. To tackle this, recent works (Niu et al., 2022, 2023; Wang et al., 2022; Shu et al., 2022) have proposed selecting confident or reliable samples for test-time model learning. This strategy significantly reduces the number of backward passes needed for an entire set of testing data, thereby enhancing adaptation efficiency and performance. Moreover, MECTA (Hong et al., 2023) proposes a series of techniques to reduce the run-time memory consumption of BP-based TTA, including reducing the batch size and stopping the BP caching heuristically. Nonetheless, these TTA methods still depend on BP, which poses challenges for resource-constrained edge devices like smartphones and FPGAs, especially those with quantized models. These devices often have limited memory and may not support backpropagation, thus hindering the broad real-world applications of BP-based TTA methods. In this context, we introduce a forward-only test-time adaptation method, aiming to achieve better performance than BP-based TTA methods but without using any backward propagation.
@@ B.2 More Evaluation Protocols
FOA (Ours). For the default configuration of hyperparameters, the number of prompt embeddings $N_{p}$ is set to 3 using uniform initialization. We use CMA-ES [^4] https://github.com/CMA-ES/pycma as the update rule, with the batch size $BS$ of 64, and the population size $K$ of $28=4+3\times\log(prompt~{}~{}dim)$ by following (Hansen, 2016). The $\lambda$ in Eqn. (3.1) is set to 0.4×$\times BS/64$ on ImageNet-C/V2/Sketch, and 0.2×$\times BS/64$ on ImageNet-R to balance the magnitude of two losses. We use the validation set of ImageNet-1K to estimate source training statistics. The moving average factor $\alpha$ in Eqn. (9) is set to 0.1. The step size $\gamma$ in Eqn. (7) is set to 1. For updating prompts and normalization layers with SGD optimizer in Table 9 with Eqn. (3.1), the entropy loss is divided by the batch size of 64 and the $\lambda$ is set to 30 to make the two losses have similar magnitude. We investigate the effects of these hyperparameters with different setups in Section 4.3 and Appendix C. The source in-distribution statistics $\{\bm{\mu}_{i}^{S},{\bm{\sigma}}_{i}^{S}\}_{i=0}^{N}$ are calculated without using the newly inserted prompt.

### v2

@@ 4.4 More Discussions
Computational Complexity Analyses. The primary computational demand of FOA stems from $K$ (the population size in CMA) forward passes. Though FOA requires more forward passes than TENT, its independence from backward passes significantly reduces memory usage and potentially boosts overall efficiency. In Table 8, our Activation Shifting achieves almost the same efficiency as NoAdapt while outperforming T3A, i.e., 59.1% vs. 56.9% accuracy. Notably, our BP-free FOA ($K\small{=}2$) matches the accuracy of BP-based TENT with lower run time and memory, and FOA ($K\small{=}6$) further surpasses BP-based CoTTA in all aspects. Moreover, even at $K\small{=}28$, FOA maintains much higher efficiency than augmentation-based methods like MEMO (Zhang et al., 2022). In FOA, $K$ is a hyperparameter that one can select different values for performance and efficiency trade-off. Note that TENT updates only norm layers, making it more efficient than full-model backpropagation.
@@
Average Run Time Memory
Method BP #FP #BP Acc. ECE (seconds) (MB)
@@
Effectiveness on ResNet (He et al., 2016) and VisionMamba (Zhu et al., 2024). For ResNet-50, we feed the original image to a learnable 7$\times$7 Conv layer to generate prompts with the same size as the image and then add prompts to the image as the model’s input. For VisionMamba, we concatenate learnable input prompts with the patch embeddings. We optimize these learnable parameters/prompts by FOA. From Table 10, FOA achieves comparable performance with gradient-based TENT and SAR on VisionMamba. While on ResNet-50, FOA outperforms BN Adapt (Schneider et al., 2020) but still suffers a large performance gap compared to TENT. This is because convolution is a local operation, making it hard to add location-invariant input prompts in CNN to affect the whole network, whereas in transformers this can be achieved using concatenation to add prompts. How to enhance the forward-only adaptation performance on ConvNets is still a promising and challenging direction, and we leave this to our future work.
Table 10: Effectiveness of FOA on ResNet and VisionMamba (Zhu et al., 2024). Results obtained on ImageNet-C (Gaussian noise, level 5). FOA is modified from FOA by replacing CMA optimizer with SGD and updating the affine parameters of norm layers.
ResNet-50 VisionMamba
Method Need BP? Acc. (%) ECE (%) Acc. (%) ECE (%)
NoAdapt
3.0 19.7 40.9 3.8
BN Adapt
16.0 1.3 n/a n/a
TENT
29.4 11.4 49.2 12.1
SAR
30.7 3.4 49.0 11.4
FOA (ours)
22.6 1.7 49.6 4.3
FOA
33.6 12.8 56.5 13.6
Effectiveness under Non-i.i.d. Scenarios. We verify the effectiveness of FOA under two non-i.i.d. scenarios by following NOTE (Gong et al., 2022) and SAR (Niu et al., 2023), i.e., online imbalanced label distribution shifts: test data come in a class order and mixed domain shifts: test data stream consists of multiple randomly mixed domains with different distribution shifts. Results in Table 11 show that FOA performs stably even under non-i.i.d. settings. In general, the performance of all methods degrades when encountering non-i.i.d. scenarios. However, FOA still achieves the best performance compared with TENT and SAR, further demonstrating the superiority of the proposed back-to-source feature alignment fitness and feature shifting scheme.
Table 11: Effectiveness of FOA under non-i.i.d. scenarios. Results obtained on ViT and ImageNet-C (level 5). For mild (i.i.d.) and online imbalanced label shift scenarios, we report the average result over 15 corruptions. For mixed shifts, the performance is evaluated on a single data stream consisting of 15 mixed corruptions.
mild scenarios online label shifts mixed shifts
Method BP? Acc. ECE Acc. ECE Acc. ECE
TENT
59.6 18.5 60.2 17.7 56.9 29.2
SAR
62.7 7.0 60.8 7.5 61.4 14.8
FOA (ours)
66.3 3.2 62.1 6.6 62.0 4.9
Comparison w.r.t. In-Distribution Performance. From Table 12, our FOA maintains almost the same in-distribution accuracy as NoAdapt, outperforming all other compared methods. This success mainly benefits from: 1) we do not modify any internal model parameters, and 2) we regularize the target test features back to the source distribution via an alignment-based fitness function and an activation shifting scheme. Both of these two components are able to alleviate the issue of catastrophic forgetting.
Table 12: Comparison w.r.t. in-distribution performance, i.e., on clean/original ImageNet validation set, with ViT as the base model.
Method NoAdapt TENT CoTTA SAR FOA (ours)
Acc. (%) 85.17 84.80(-0.37) 83.91(-1.26) 84.52(-0.65) 85.11(-0.06)
ECE (%) 9.6 3.9 1.8 0.9 2.1
Differences from Previous Forward-Only TTA. The main difference is that we explicitly exploit model feedback to facilitate optimization with learnable parameters, while prior forward-only TTA methods do not. For example, BN statistics calibration-based methods, including SITA (Khurana et al., 2021) exploit test data to calculate the mean and variances in batch norm layers, T3A (Iwasawa & Matsuo, 2021) adjusts class prototypes for adaptation, DDA (Gao et al., 2023) conducts diffusion on input images to remove the potential distribution shifts, etc. However, these methods do not actively learn from the model’s feedback w.r.t. a given test sample, and thus their performance on OOD data may be limited. In contrast, our FOA conducts explicit input prompt learning according to the model’s feedback, i.e., using the proposed fitness function consisting of prediction entropy and feature discrepancy, thereby achieving much better adaptation performance.
@@ 5 Conclusion
In this paper, we aim to implement online test-time adaptation without the need for backpropagation and altering the model parameters. This advancement highly broadens the scope of TTA’s real-world applications, particularly in resource-limited scenarios such as smartphones and FPGAs where backpropagation is often not feasible. To this end, we propose a test-time Forward-Optimization Adaptation (FOA) method. In FOA, we online learn an input prompt through a covariance matrix adaptation technique, paired with a designed unsupervised fitness function to provide stable learning signals. Moreover, we devise a “back-to-source” activation shifting scheme to directly alter the activations from the out-of-distribution domain to the source in-distribution domain, further boosting the adaptation performance. Extensive experiments on four large-scale out-of-distribution benchmarks with full precision (32-bit) and quantized (8-bit/6-bit) models verify our superiority.
Acknowledgment
This research was supported, in part, by the Joint NTU-WeBank Research Centre on Fintech, Nanyang Technological University, Singapore, and the National Research Foundation, Singapore under its Industry Alignment Fund – Pre-positioning (IAF-PP) Funding Initiative. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of National Research Foundation, Singapore. The computational work was partially performed on resources of the National Supercomputing Centre, Singapore.
Impact Statement
@@ Appendix A Related Work
$\bullet$ Backpropagation (BP)-Free TTA. In the early development of BP-free TTA, attention was primarily given to adjusting batch normalization (BN) layer statistics by computing mean and variance from testing data (Nado et al., 2020; Schneider et al., 2020; Gong et al., 2022). This method, however, assumes multiple test samples for each prediction. To address this, later studies proposed single-sample BN adaptation techniques, such as using data augmentation (Khurana et al., 2021), mix-up training and testing statistics (Hu et al., 2021; Lim et al., 2023). Moreover, NOTE (Gong et al., 2022) proposes instance-aware batch normalization and DELTA (Zhao et al., 2023) exploits test-time batch re-normalization to adapt a given model under non-i.i.d. testing scenarios. In addition to BN adaptation, other methodologies have been explored and shown to be effective, such as prototype-based classifier adjustment (Iwasawa & Matsuo, 2021), predicted logits correction (Boudiaf et al., 2022), diffusion-based input image adaptation (Gao et al., 2023; Oh et al., 2024). However, since BP-free TTA does not update the core model parameters, it might exhibit limited learning capabilities, often resulting in suboptimal performance when dealing with out-of-distribution testing data. Therefore, BP-based TTA emerged as an effective solution to significantly boost the out-of-distribution generalization performance.
@@
$\bullet$ Backpropagation-Based TTA. One pioneering work of BP-based TTA is known as Test-Time Training (TTT) (Sun et al., 2020). For TTT methods, they first train a source model using both supervised and self-supervised objectives, followed by adapting the model at test time with the self-supervised objective, such as rotation prediction (Sun et al., 2020), contrastive learning (Liu et al., 2021a; Bartler et al., 2022), reconstruction learning (Gandelsman et al., 2022; Deng et al., 2024). To avoid directly altering the model training process and access to source data, Fully TTA methods update any given model via unsupervised learning objectives, such as entropy minimization (Wang et al., 2021; Wen et al., 2023; Niu et al., 2023; Tan et al., 2024), prediction consistency maximization (Zhang et al., 2022; Fleuret et al., 2021; Zeng et al., 2023) and feature distribution alignment (Mirza et al., 2023; Lin et al., 2023).
@@
However, BP-based TTA methods typically require multiple backward propagations for each test sample, leading to computational inefficiency. To tackle this, recent works (Niu et al., 2022, 2023; Wang et al., 2022; Shu et al., 2022) have proposed selecting confident or reliable samples for test-time model learning. This strategy significantly reduces the number of backward passes needed for an entire set of testing data, thereby enhancing adaptation efficiency and performance. Moreover, MECTA (Hong et al., 2023) proposes a series of techniques to reduce the run-time memory consumption of BP-based TTA, including reducing the batch size and stopping the BP caching heuristically. Nonetheless, these TTA methods still depend on BP, which poses challenges for resource-constrained edge devices like smartphones and FPGAs, especially those with quantized models. These devices often have limited memory and may not support backpropagation, thus hindering the broad real-world applications of BP-based TTA methods. In this context, we introduce a forward-only and optimization-based TTA method, aiming to achieve better performance than BP-based TTA methods but without using any backward propagation.
@@
Connection to Conventional Alignment-based Domain Adaptation Methods. Conventional DA methods often seek to learn domain-invariant features by minimizing the Maximum Mean Discrepancy (MMD) between the source and target distributions (Long et al., 2015, 2016; Abuduweili et al., 2021). BUFR (Eastwood et al., 2022) extends this idea to a source-free manner and directly minimizes the KL divergence between the source and target feature distributions. Our proposed Activation Shifting scheme follows the general idea of source-target alignment but differs from them mainly in two aspects. 1) Conventional DA methods often calculate the MMD or KL divergence as a differentiable loss function and exploit it for gradient backpropagation, thereby updating model weights. In contrast, our Activation Shifting directly edits the activations back to the source distribution in a gradient-free manner without modifying model weights, which facilitates efficient test-time adaptation. 2) Previous DA methods typically work offline, i.e., training the model on a pre-collected dataset for several epochs. Conversely, our Activation Shifting approach is effective for online test data streams, where adaptation occurs immediately after each mini-batch of test samples is processed (e.g., with batch sizes of 2 or 4).
@@ B.2 More Evaluation Protocols
FOA (Ours). For the default configuration of hyperparameters, the number of prompt embeddings $N_{p}$ is set to 3 using uniform initialization. We use CMA-ES [^4] https://github.com/CMA-ES/pycma as the update rule, with the batch size $BS$ of 64, and the population size $K$ of $28=4+3\times\log(prompt~{}~{}dim)$ by following (Hansen, 2016). The $\lambda$ in Eqn. (3.1) is set to 0.4×$\times BS/64$ on ImageNet-C/V2/Sketch, and 0.2×$\times BS/64$ on ImageNet-R to balance the magnitude of two losses. We use the validation set of ImageNet-1K to estimate source training statistics. The moving average factor $\alpha$ in Eqn. (9) is set to 0.1. The step size $\gamma$ in Eqn. (7) is set to 1.0, aiming to exactly align the overall center of testing and training features. For updating prompts and normalization layers with SGD optimizer in Table 9 with Eqn. (3.1), the entropy loss is divided by the batch size of 64 and the $\lambda$ is set to 30 to make the two losses have similar magnitude. We investigate the effects of these hyperparameters with different setups in Section 4.3 and Appendix C. The source in-distribution statistics $\{\bm{\mu}_{i}^{S},{\bm{\sigma}}_{i}^{S}\}_{i=0}^{N}$ are calculated without using the newly inserted prompt.
@@ Appendix C More Ablation Studies
Comparison with MEMO (Zhang et al., 2022) under Different Number of Test Samples. To verify the effectiveness of FOA when only limited test samples are available for adaptation, we record the model’s accuracy on the entire test set after the online adaption on $N$ test samples. From Figure 4, our FOA outperforms MEMO and NoAdapt at the beginning of the test data stream, i.e., less than 200 samples, showing that FOA does not rely on a long test data stream. The good performance mainly benefits from our Activation Shifting scheme, which boosts the performance a lot in cases of the CMA-based prompt adaptation is inadequate, e.g., at the beginning of adaptation. Here, FOA with a small batch size (BS) of 4 adapts faster, as a smaller batch size leads to more learning steps.
Figure 4: Online accuracy comparison with MEMO (Zhang et al., 2022) on ViT and ImageNet-C (Gaussian noise, severity level 5).